OpenStudy (anonymous):

14. The perimeter of a rectangle is 62. The area of the same rectangle is 234. Which of the following equations could be used to find the dimensions of the rectangle? A. w(31 – w) = 234 B. w(62 – w) = 234 C. w(31 + w) = 234 D. w(62 + w) = 234

OpenStudy (anonymous):

Please help!

OpenStudy (mertsj):

how do you find perimeter of a rectangle?

OpenStudy (anonymous):

P=2(l+w)

OpenStudy (mertsj):

And how do you find the area or a rectangle?

OpenStudy (anonymous):

A=lw

OpenStudy (mertsj):

So for this particular problem, we could write: 2l+2w=62 and lw = 234. Would you agree?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

But then, looking at the answer choices, I see that they must be looking for an expression for the area because they are all equal to 234 which is the area. Do you agree?

OpenStudy (anonymous):

Yes, I agree

OpenStudy (mertsj):

And furthermore, I see that the length has been eliminated somehow because the only variable in the area statement is w. So I need a replacement for the "l" in A=lw. And the only place I have to obtain such a replacement is from the perimeter equation.

OpenStudy (mertsj):

Does this still make sense?

OpenStudy (anonymous):

Yes, yes it does

OpenStudy (mertsj):

So it looks like we had better take our perimeter equation: 2l+2w = 62 and solve it for l and put that value in the area equation: lw = 234 in place of l and see what we get.

OpenStudy (anonymous):

Ok then. thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Would the right answer be B? @Mertsj

OpenStudy (mertsj):

no. When you divide by 2, make sure you divide EVERY term by 2

OpenStudy (anonymous):

Is A the correct answer?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Thank you so much!!!:)

OpenStudy (mertsj):

yw